OpenStudy (kittiwitti1):

Several questions, so if you have time please stay tuned. If not, just answer the one that you see below 1. A and B are points on the graphs of functions y1 and y2, respectively. Find the corresponding point on the graph of y = y1 + y2. A = (3, 3), B = (3, 2)

OpenStudy (dayakar):

u there

OpenStudy (dayakar):

where is graph

OpenStudy (kittiwitti1):

graph?

OpenStudy (kittiwitti1):

there is no graph http://icecream.me/1036ae2a33e54e2a4d4dcffe24a1282a

OpenStudy (dayakar):

can u identify y1,y2

OpenStudy (kittiwitti1):

um

OpenStudy (dayakar):

A=(x1,y1)=(3,3) B=(x2,y2)=(3,2)

OpenStudy (kittiwitti1):

yes. sorry, internet lagged

OpenStudy (dayakar):

find y = y1+y2

OpenStudy (kittiwitti1):

3+2=5

OpenStudy (dayakar):

,very good

OpenStudy (kittiwitti1):

well. that was easier than I thought it would be, thought it would be more complex of a solution

OpenStudy (dayakar):

in two points compare x value

OpenStudy (kittiwitti1):

ok, next one http://icecream.me/cf65b9e68c8062226ed7d4b061c81bcc

OpenStudy (dayakar):

it was not completed

OpenStudy (kittiwitti1):

?

OpenStudy (kittiwitti1):

I put 3,5 and got right answer

OpenStudy (kittiwitti1):

because only the y value changes yes?

OpenStudy (dayakar):

(x,y)=(3,5) excellent

OpenStudy (kittiwitti1):

okay. what about the next one ? :o

OpenStudy (dayakar):

The arccosine of x is defined as the inverse cosine function of x when -1≤x≤1.

OpenStudy (kittiwitti1):

oh

OpenStudy (kittiwitti1):

well yes, but that isn't an option from what I saw o-o

OpenStudy (dayakar):

it is the third one

OpenStudy (kittiwitti1):

o-o

OpenStudy (kittiwitti1):

Wait what?! D:

OpenStudy (dayakar):

C

OpenStudy (kittiwitti1):

I considered, but I don't see how x is the angle

OpenStudy (kittiwitti1):

...nvm it was #4

OpenStudy (kittiwitti1):

http://icecream.me/e7f2c88eeeae2929c3758ac63242439e

OpenStudy (dayakar):

u r right

OpenStudy (dayakar):

any more

OpenStudy (kittiwitti1):

Hmm...

OpenStudy (dayakar):

mean

OpenStudy (dayakar):

may i leave

OpenStudy (kittiwitti1):

@dayakar what? my internet lagged

OpenStudy (kittiwitti1):

here is another one Fill in each blank with the appropriate number. (Select all that apply.) If y = sin−1 x, then y can terminate in quadrant ____ or ____.

OpenStudy (dayakar):

I and III

OpenStudy (dayakar):

any more

OpenStudy (kittiwitti1):

whoa when did you come back did my internet lag again =_=

OpenStudy (kittiwitti1):

@dayakar it was wrong...

OpenStudy (kittiwitti1):

@Directrix

Directrix (directrix):

What is the question?

OpenStudy (kittiwitti1):

Fill in each blank with the appropriate number. (Select all that apply.) If y = sin−1 x, then y can terminate in quadrant ____ or ____.

Directrix (directrix):

Based on the attachment, Quads I and IV. http://www.themathpage.com/atrig/inversetrig.htm

OpenStudy (kittiwitti1):

Oh, so it WAS QIV lol

OpenStudy (kittiwitti1):

I'd been debating that but my brain isn't processing things correctly. x_x

OpenStudy (kittiwitti1):

If y = cos−1 x, then y can terminate in quadrant ____ or ____. Then this would be QI and Q.. 3?

Directrix (directrix):

>>Then this would be QI and Q.. 3? No. See attachment.

OpenStudy (kittiwitti1):

Ergh, Q2. Sorry

OpenStudy (kittiwitti1):

I thought I had been wrong, so I wrote 3. e_e

OpenStudy (kittiwitti1):

http://icecream.me/7f8680e342b20f2e81460c7b94770674

Directrix (directrix):

Please begin a new thread. It is customary here to post one question per thread. Thanks.

OpenStudy (kittiwitti1):

Oh... but I did say there were several questions to be asked

OpenStudy (kittiwitti1):

I also like to keep my questions asked:questions answered ratio low...